# FedPeWS: Personalized Warmup via Subnetworks for Enhanced Heterogeneous Federated Learning

### v1

@@ 5.1 Datasets and Network Architecture
Figure 3: Samples from the custom synthetic dataset.
@@
We integrate two distinct datasets, CIFAR-10 [22] and MNIST [23], to explore how different clients might adapt when faced with disparate data sources. CIFAR-10 comprises of $32\times 32$ pixel images categorized into 10 object classes. MNIST, typically featuring $28\times 28$ pixel images across 10 digit classes, is upscaled to $32\times 32$ pixel to standardize dimensions with CIFAR-10. We compile a balanced dataset by randomly selecting 400 samples from each class for the training set and 200 samples for the test set from the combined pool of 20 classes. This setup aims to simulate a FL environment where multiple clients handle significantly varied data types. For this dataset, we employ a convolutional neural network comprising four convolutional layers, each having a kernel size of 3 and padding of 1, followed by max pooling. This is succeeded by a fully connected layer. This architecture was used because of its simplicity and widespread use in the literature [50, 15].
@@
We amalgamate three distinct medical datasets: PathMNIST, OCTMNIST, and TissueMNIST [48], to develop a universal medical prognosis model capable of recognizing various tasks using a single model. The datasets contain 9, 4, and 8 classes, respectively, totaling to 21 classes. For this dataset, we utilized the same architecture and training details described in the CIFAR-MNIST dataset.
@@ 5.2 Experimental Setup
For scenarios with a smaller number of collaborators ($N={2,3,4}$), we manually partition the training dataset to tailor the data distribution to specific participants. In the $N=2$ scenario, we partition as follows: (i) For the Synthetic dataset, encompassing both Synthetic-32K and Synthetic-3.2K, even-numbered classes are assigned to participant 1, while odd-numbered classes are allocated to participant 2. (ii) For the CIFAR-MNIST combination, all CIFAR-10 samples are assigned to participant 1, with MNIST samples allocated to participant 2. In the $N=3$ scenario, the {Path-OCT-Tissue}MNIST dataset is partitioned into three splits corresponding to the individual datasets, with PathMNIST assigned to participant 1, OCTMNIST to participant 2, and TissueMNIST to participant 3. For the $N=4$ scenario, the synthetic dataset is divided so that each class is exclusively allocated to one of the four participants.
For scenarios with a larger number of participants ($N\geq 10$), we employ a Dirichlet distribution to partition the training set. This approach utilizes a concentration parameter $\alpha$ to simulate both homogeneous and heterogeneous data distributions [51, 27, 32, 44]. We experiment with various values of $\alpha$, specifically $\alpha\in\{0.1,0.5,1.0,2.0,5.0\}$, to explore the effects of dataset heterogeneity (lower $\alpha$ values) and homogeneity (higher $\alpha$ values) on the model performance. This methodological diversity allows us to comprehensively assess our approach under varying data conditions. Results for large $N$ ($>100$) are reported in the appendix.
@@
In federated optimization, we primarily benchmark against the FedAvg algorithm [33], a standard approach in federated learning. However, our algorithm is designed to be versatile, functioning as a ‘plug-and-play’ solution that is compatible with various other optimizers. To demonstrate this adaptability, we also conduct experiments using FedProx [29], showcasing our method’s capabilities across different optimization frameworks. For our experiments, we fix the local learning rate $\eta_{\ell}=0.001$ in the Synthetic-32K dataset case, and we set $\eta_{\ell}=0.01$ for other experiments. Also, the mask learning rate is fixed $\eta_{s}=0.1$. Furthermore, we vary the global learning rate $\eta_{g}\in\{0.1,0.25,0.5,1.0\}$ to observe the differences in optimization behavior between the baseline and our proposed methods. Additionally, we employ two distinct batch sizes $\{32,8\}$ for Synthetic-32K and Synthetic-3.2K, respectively. For experiments involving the CIFAR-MNIST and {Path-OCT-Tissue}MNIST datasets, we standardize the batch size to 64. We conduct our experiments on NVIDIA RTX A6000 GPUs on an internal cluster server, with each run utilizing a single GPU. The execution time for each run is capped at less than an hour, which indicates the maximum execution time rather than the average. All results are averaged over three independent runs and the average accuracy is reported on the global test dataset.
@@
FedPeWS $\mathbf{115\pm 7.21}$ $\mathbf{182\pm 6.81}$ $\mathbf{286\pm 7.93}$ $\mathbf{301\pm 10.59}$
@@
FedPeWS $\mathbf{99.96\pm 0.01}$ $\mathbf{99.49\pm 0.60}$ $\mathbf{83.50\pm 3.52}$ $\mathbf{99.66\pm 0.19}$
Figure 4: Results of the experiments on Synthetic-{32, 3.2}K datasets with batch sizes {32, 8}, with different global learning rates $\eta_{g}\in\{1.0,0.5,0.25,0.1\}$ and communication rounds $T\in\{200,250,400,500\}$. Refer to Table 1 for the corresponding numbers. In all the above scenarios, FedPeWS converges faster to a better solution compared to FedAvg.
@@ 5.3 Experimental Results
Figure 5: Visualization of validation accuracy and loss on the Synthetic-32K dataset with $N=4$ participants and a global learning rate $\eta_{g}=1.0$.
Our experimental analysis focuses on assessing the performance of our proposed FedPeWS algorithm within the FL framework. The key findings from our studies are as follows: (i) The FedPeWS approach demonstrates a significant reduction in the number of communication rounds required to achieve target accuracy while also enhancing the final accuracy post-convergence. (ii) The FedPeWS algorithm is robust across different levels of data heterogeneity. (iii) In scenarios where full knowledge of the participant data distributions is available, the server can employ the FedPeWS-Fixed method (Figure 6). While the FedPeWS-Fixed variant shows competitive effectiveness comparable to our primary FedPeWS algorithm, the latter offers broader applicability in real-world settings.
@@
We initially report the required number of communication rounds to reach the target accuracy and the final accuracy after $T$ communication rounds for the synthetic dataset in Table 1. The results underscore that the incorporation of a personalized warmup phase in a federated setup significantly reduces the required number of communication rounds across all tested scenarios. Notably, in specific instances, such as with the Synthetic-32K dataset and $\eta_{g}=0.25$, the conventional FedAvg algorithm does not meet the target accuracy within the $T$ communication rounds. Conversely, in scenarios where $\eta_{g}\in\{0.25,0.1\}$, FedAvg only achieves the target accuracy in one of the seeds, exhibiting suboptimal performance in the other two runs. From Figure 4, it is evident that our proposed FedPeWS algorithm surpasses FedAvg in both communication efficiency and accuracy.
We also consider a more extreme data heterogeneity scenario with $N=4$ participants, depicted in Figure 5, where FedAvg completely fails by reaching only $58.4\pm 2.33\%$, whereas our FedPeWS approach reaches $91.13\pm 3.55\%$ accuracy by significantly outperforming the base optimizer (FedAvg) with a gain of $\mathbf{32.72\%}$. It is crucial to highlight that in this experiment, we set $\lambda=0.0$, effectively not enforcing diversity as outlined in Equation 6. This approach focuses solely on optimizing the masks using the first loss component, which depends only on the data distributions of each participant. This shows that, in specific scenarios, we can learn the personalized masks (Procedure I) without the need to adjust the $\lambda$ parameter, while still achieving a better performance than the base optimizer.
@@
Figure 6: Results for experiments on (a) the CIFAR-MNIST and (b) {Path-OCT-Tissue}MNIST datasets with a communication budget of $T=300$. (a) Left: Participant 1 has MNIST data samples; Participant 2 has CIFAR-10 data samples. (a) Right: Ablation study for $\lambda$ and $\tau$ parameters on CIFAR-MNIST (see Table 6). (b) Left: Each of $N=3$ participants holds unique dataset samples from {PathMNIST, OCTMNIST, TissueMNIST} pool. (b) Right: Ablation study for $\lambda$ and $\tau$ on the respective dataset (see Table 7). The first column ($\tau=0.0$) corresponds to the FedAvg algorithm. The last row presents results for the FedPeWS-Fixed algorithm.
@@
Figure 6 showcases the results of experiments on the CIFAR-MNIST dataset with $N=2$ participants and {Path-OCT-Tissue}MNIST dataset with $N=3$ participants. The left-side plots of Figures 6(a) and 6(b), which depict the performance of the global model (averaged over 3 runs), demonstrate that our method consistently achieves higher accuracy. The right side figures feature heatmap plots that annotate the global model accuracy obtained varying $\lambda\in\{0,0.1,0.3,0.5,1,2,5,10,100,1000\}$ and $\tau\in\{0.0,0.2,0.4,0.5,0.6,0.8\}$ parameters. An additional row labeled $(\lambda=-)$ represents the FedPeWS-Fixed approach, where user(server)-defined fixed masks are employed. In this method, we simply split the full network into $N$ partitions, with each partition assigned to a participant (for detailed instructions on setting masks, please see Section B.2 in the appendix). The results indicate that our approach has a low sensitivity to variations in $\lambda$ and $\tau$. For more detailed insights, please refer to Tables 6 and 7 in the appendix.
Varying degrees of heterogeneity.
[…]
Dataset /// Batch size Synthetic-32K, $32$ Synthetic-3.2K, $8$
Target accuracy $\upsilon(\%)$ $99$ $90$ $75$ $99$
No. of rounds to
reach target accuracy
FedProx $138\pm 13.22$$\times$ $\times$ $362\pm 20.00$
FedPeWS $\mathbf{115\pm 5.29}$ $\mathbf{211\pm 16.52}$ $\mathbf{314\pm 27.83}$ $\mathbf{344\pm 27.30}$
Final accuracy after
$T$ collaboration rounds
FedPeWS $\mathbf{99.96\pm 0.01}$ $\mathbf{98.40\pm 1.84}$ $\mathbf{90.40\pm 3.91}$ $\mathbf{99.92\pm 0.07}$
FedProx.
@@ 6 Conclusion
In this work, we introduced a novel concept called personalized warmup via subnetworks for heterogeneous FL -- a strategy that enhances convergence speed and can seamlessly integrate with existing optimization techniques. Results demonstrate that the proposed FedPeWS approach and achieves higher accuracy than the relevant baselines, especially when there is extreme statistical heterogeneity. Limitations of FedPeWS include the need to tune two additional hyperparameters (no. of warmup rounds and mask diversity weight) and the lack of theoretical convergence analysis.
Appendix A Related Work
Table 3 shows the comparison of our proposed approach to the existing literature.
Shared
Global Model
Level of Mask
Personalization
Learnable
Mask
Learnable
Parameters
FedPM [15] parameter-level $\bm{\times}$
IST [50] neuron-level $\bm{\times}$ (random)
LTH [10] $\bm{\times}$ parameter-level $\bm{\times}$ (pruned)
FedWeIT [49] $\bm{\times}$ parameter-level
FjORD [14] parameter-level $\bm{\times}$ (slimmed)
FedPeWS (ours) neuron-level
@@ B.2 Fixed Mask Generation
Figure 9 illustrates how we design masks for FedPeWS-Fixed experiments in scenarios with $N=2$ participants. For cases involving $N=4$ participants, the full network $\mathcal{M}_{x}$ (classifier $\mathcal{M}$ parameterized with $x$) is divided into four subnetworks, vertically, with each subnetwork corresponding to one of the participants. As such, we vertically partition the hidden neurons in the network into $N$ groups (subnetworks) and design the mask to assign each group to one participant, ensuring no overlap. This design choice is based on the assumption that classes held by each participant are highly heterogeneous, thus preventing any intersection in the masks. This setting is specifically tailored for the FedPeWS-Fixed method and doesn’t necessitate performing optimization over the masks $m_{i}$, they are kept fixed.
@@ Appendix C Experimental Results
C.1 Wall-clock Time vs. Accuracy
Figure 10 illustrates the wall-clock time versus accuracy results, which correspond to Figure 4 in the main paper. From this comparison, FedPeWS demonstrates a slightly improved performance over FedAvg in terms of wall-clock time in two of the four scenarios. However, it underperforms slightly in the remaining two scenarios, with only a marginal increase in time. This variance is attributed to the alternation between training masks and weights during the warm-up phase, impacting the time efficiency.
Figure 10: Wall-clock time vs. accuracy plot corresponding to Figure 4 of the main paper.
C.2 FedPeWS-Fixed. Mask Length Study

### v2

@@ 5.1 Datasets and Network Architecture
Figure 3: Samples from the custom synthetic dataset.
@@
We integrate two distinct datasets, CIFAR-10 [50] and MNIST [51], to explore how different clients might adapt when faced with disparate data sources. CIFAR-10 comprises $32\times 32$ pixel images categorized into 10 object classes. MNIST, typically featuring $28\times 28$ pixel images across 10 digit classes, is upscaled to $32\times 32$ pixel to standardize dimensions with CIFAR-10. We compile a balanced dataset by randomly selecting 400 samples from each class for the training set and 200 samples for the test set from the combined pool of 20 classes. This setup aims to simulate a FL environment where multiple clients handle significantly varied data types. For this dataset, we employ a convolutional neural network comprising four convolutional layers, each having a kernel size of 3 and padding of 1, followed by max pooling. This is succeeded by a fully connected layer. This architecture was used because of its simplicity and widespread use in the literature [35, 20].
@@
We amalgamate three distinct medical datasets: PathMNIST, OCTMNIST, and TissueMNIST [52], to develop a universal medical prognosis model capable of recognizing various tasks using a single model. The datasets contain 9, 4, and 8 classes, respectively, totaling 21 classes. For this dataset, we utilized the same architecture and training details described in the CIFAR-MNIST dataset.
@@ 5.2 Experimental Setup
For scenarios with a smaller number of collaborators ($N={2,3,4}$), we manually partition the training dataset to tailor the data distribution to specific participants. In the $N=2$ scenario, we partition as follows: (i) For the Synthetic dataset, encompassing both Synthetic-32K and Synthetic-3.2K, even-numbered classes are assigned to participant 1, while odd-numbered classes are allocated to participant 2. (ii) For the CIFAR-MNIST combination, all CIFAR-10 samples are assigned to participant 1, with MNIST samples allocated to participant 2. In the $N=3$ scenario, the {Path-OCT-Tissue}MNIST dataset is partitioned into three splits corresponding to the individual datasets, with PathMNIST assigned to participant 1, OCTMNIST to participant 2, and TissueMNIST to participant 3. For the $N=4$ scenario, the synthetic dataset is divided so that each class is exclusively allocated to one of the four participants. For scenarios with a larger number of participants ($N\geq 10$), we employ a Dirichlet distribution to partition the training set. This approach utilizes a concentration parameter $\alpha$ to simulate both homogeneous and heterogeneous data distributions [53, 54, 55, 28]. We experiment with various values of $\alpha$, specifically $\alpha\in\{0.1,0.5,1.0,2.0,5.0\}$, to explore the effects of dataset heterogeneity (lower $\alpha$ values) and homogeneity (higher $\alpha$ values) on the model performance. This methodological diversity allows us to comprehensively assess our approach under varying data conditions. Results for large $N$ ($>100$) are reported in the appendix.
@@
In federated optimization, we primarily benchmark against the FedAvg algorithm [1], a standard approach in federated learning. However, our algorithm is designed to be versatile, functioning as a ‘plug-and-play’ solution that is compatible with various other optimizers. To demonstrate this adaptability, we also conduct experiments using FedProx [6], showcasing our method’s capabilities across different optimization frameworks. Moreover, we compare these two variants to SCAFFOLD [7], FedNova [27], and MOON [56]. For our experiments, we fix the local learning rate $\eta_{\ell}=0.001$ in the Synthetic-32K dataset case, and we set $\eta_{\ell}=0.01$ for other experiments. Also, the mask learning rate is fixed $\eta_{s}=0.1$. Furthermore, we vary the global learning rate $\eta_{g}\in\{0.1,0.25,0.5,1.0\}$ to observe the differences in optimization behavior between the baseline and our proposed methods. Additionally, we employ two distinct batch sizes $\{32,8\}$ for Synthetic-32K and Synthetic-3.2K, respectively. For experiments involving the CIFAR-MNIST and {Path-OCT-Tissue}MNIST datasets, we standardize the batch size to 64. We conduct our experiments on NVIDIA RTX A6000 GPUs on an internal cluster server, with each run utilizing a single GPU. The execution time for each run is capped at less than an hour, which indicates the maximum execution time rather than the average. All results are averaged over three independent runs and the average accuracy is reported on the global test dataset.
@@
FedAvg+PeWS $\mathbf{115\pm 7.21}$ $\mathbf{182\pm 6.81}$ $\mathbf{286\pm 7.93}$ $\mathbf{301\pm 10.59}$
@@
FedAvg+PeWS $\mathbf{99.96\pm 0.01}$ $\mathbf{99.49\pm 0.60}$ $\mathbf{83.50\pm 3.52}$ $\mathbf{99.66\pm 0.19}$
Figure 4: Results on Synthetic-{32, 3.2}K datasets with batch sizes {32, 8}, global learning rates $\eta_{g}\in\{1.0,0.5,0.25,0.1\}$ and communication rounds $T\in\{200,250,400,500\}$. See Table 2 for details. FedPeWS consistently converges faster and outperforms FedAvg.
@@ 5.3 Experimental Results
Our experimental analysis focuses on assessing the performance of our proposed FedPeWS algorithm within the FL framework. The key findings from our studies are as follows: (i) The FedPeWS approach demonstrates a significant reduction in the number of communication rounds required to achieve target accuracy while also enhancing the final accuracy post-convergence. (ii) The FedPeWS algorithm is robust across different levels of data heterogeneity. (iii) In scenarios where full knowledge of the participant data distributions is available, the server can employ FedPeWS-Fixed method (Figure 6).
@@
Figure 5: Visualization of validation accuracy and loss on the Synthetic-32K dataset with $N=4$.
We initially report the required number of communication rounds to reach the target accuracy and the final accuracy after $T$ communication rounds for the synthetic dataset in Table 2. The results underscore that the incorporation of a personalized warmup phase in a federated setup significantly reduces the required number of communication rounds across all tested scenarios. Notably, in specific instances, such as with the Synthetic-32K dataset and $\eta_{g}=0.25$, the conventional FedAvg algorithm does not meet the target accuracy within the $T$ communication rounds. Conversely, in scenarios where $\eta_{g}\in\{0.25,0.1\}$, FedAvg only achieves the target accuracy in one of the seeds, exhibiting suboptimal performance in the other two runs. From Figure 4, it is evident that our proposed FedPeWS algorithm surpasses FedAvg in both communication efficiency and accuracy. We also consider a more extreme data heterogeneity scenario with $N=4$ participants, depicted in Figure 5, where FedAvg completely fails by reaching only $58.4\pm 2.33\%$, whereas our FedPeWS approach reaches $91.13\pm 3.55\%$ accuracy by significantly outperforming the base optimizer (FedAvg) with a gain of $\mathbf{32.72\%}$. It is crucial to highlight that in this experiment, we set $\lambda=0.0$, effectively not enforcing diversity as outlined in Equation 6. This approach focuses solely on optimizing the masks using the first loss component, which depends only on the data distributions of each participant. This shows that, in specific scenarios, we can learn the personalized masks (Procedure I) without the need to adjust the $\lambda$ parameter, while still achieving a better performance than the base optimizer.
@@
Figure 6: Results for experiments with $T=300$ on (a) CIFAR-MNIST and (b) {Path-OCT-Tissue}MNIST datasets. (a) Participant 1 uses MNIST; Participant 2 uses CIFAR-10; ablation study for $\lambda$ and $\tau$ (Table 7). (b) $N=3$ participants use {PathMNIST, OCTMNIST, TissueMNIST}; ablation study for $\lambda$ and $\tau$ (Table 8). FedPeWS-Fixed results appear in the last row; $\tau=0.0$ denotes FedAvg.
@@
Figure 6 showcases the results of experiments on the CIFAR-MNIST dataset with $N=2$ participants and {Path-OCT-Tissue}MNIST dataset with $N=3$ participants. The left-side plots of Figures 6(a) and 6(b), which depict the performance of the global model (averaged over 3 runs), demonstrate that our method consistently achieves higher accuracy. The right side figures feature heatmap plots that annotate the global model accuracy obtained varying $\lambda\in[0,1000]$ and $\tau\in[0.0,0.8]$ parameters. An additional row $(\lambda=-)$ represents the FedPeWS-Fixed approach, where user(server)-defined fixed masks are employed. In this method, we simply split the full network into $N$ partitions, with each partition assigned to a participant (for detailed instructions see Section B.2 in the appendix). The results indicate that our approach has a low sensitivity to variations in $\lambda$ and $\tau$. For more detailed insights, please refer to Tables 7 and 8 in the appendix.
Table 3: Comparison to the SOTA algorithms.
Dataset CIFAR-MNIST {P-O-T}MNIST
Fedavg $71.78\pm 0.66$ $52.83\pm 1.26$
FedProx $72.27\pm 0.88$ $51.28\pm 1.03$
SCAFFOLD $71.83\pm 0.24$ $53.05\pm 0.60$
FedNova $71.63\pm 0.98$ $53.05\pm 0.83$
MOON $71.84\pm 1.09$ $52.10\pm 0.19$
FedAvg+PeWS $\mathbf{75.83\pm 0.88}$ $\mathbf{55.12\pm 0.56}$
FedProx+PeWS $75.04\pm 0.85$ $54.67\pm 0.43$
Comparison to the state-of-the-art (SOTA) algorithms.
We evaluate our proposed algorithm, FedPeWS, against state-of-the-art (SOTA) algorithms, with the results summarized in Table 3. As highlighted in the table, algorithms such as FedProx, SCAFFOLD, FedNova, and MOON demonstrate performance comparable to FedAvg, as illustrated in Figure 6. In contrast, our algorithm, when applied on top of FedAvg and FedProx, achieves superior performance.
@@ 6 Conclusion
In this work, we introduced a novel concept called personalized warmup via subnetworks for heterogeneous FL – a strategy that enhances convergence speed and can seamlessly integrate with existing optimization techniques. Results demonstrate that the proposed FedPeWS approach achieves higher accuracy than the relevant baselines, especially when there is extreme statistical heterogeneity.
Acknowledgements
This material is partly based on work supported by the Office of Naval Research N00014-24-1-2168.
Appendix A Scalability of the Mask Mechanism
Our approach is designed as a plug-and-play extension to existing federated optimization methods with minimal overhead.
To estimate the computational overhead, let us assume that a baseline method such as FedAvg incurs a training cost of 1 unit per communication round. In our method, a warm-up phase occupies a fraction $\tau$ of the total communication rounds (e.g., $\tau\in\{0.1,0.5\}$). Let $\omega$ be the computational complexity of each warmup round. Consequently, the total cost of the overall training will increase only when $\omega$ is greater than 1. During the warmup rounds, additional costs from mask learning are incurred. It must be noted that we learn only neuron-level masks, and the number of hidden neurons is much lower than the number of parameters. On the other hand, these additional mask learning costs are offset by the fact that only sparse subnetworks are learned during the warmup rounds. Empirically, the observed sparsification level lies in the range $[0.1,0.5]$. Under the worst-case scenario of sparsification at 0.5, these two factors (computational cost increase due to mask learning and computational cost decrease due to sparsification) cancel each other out and $\omega\approx 1$. Therefore, the computational cost of FedPeWS is approximately the same as FedAvg when sparsification is $0.5$. For more aggressive sparsification ($<0.5$), FedPeWS becomes more computationally efficient than FedAvg.
@@ B.2 Fixed Mask Generation
Figure 7 illustrates how we design masks for FedPeWS-Fixed experiments in scenarios with $N=2$ participants. For cases involving $N=4$ participants, the full network $\mathcal{M}_{x}$ (classifier $\mathcal{M}$ parameterized with $x$) is divided into four subnetworks, vertically, with each subnetwork corresponding to one of the participants. As such, we vertically partition the hidden neurons in the network into $N$ groups (subnetworks) and design the mask to assign each group to one participant, ensuring no overlap. This design choice is based on the assumption that classes held by each participant are highly heterogeneous, thus preventing any intersection in the masks. This setting is specifically tailored for the FedPeWS-Fixed method and doesn’t necessitate performing optimization over the masks $m_{i}$; they are kept fixed.
@@ Appendix C Experimental Results
C.1 Varying Degrees of Heterogeneity
C.2 Comparison to FedProx
[…]
Dataset /// Batch size Synthetic-32K, $32$ Synthetic-3.2K, $8$
Target accuracy $\upsilon(\%)$ $99$ $90$ $75$ $99$
No. of rounds to
reach target accuracy
FedProx $138\pm 13.22$$\times$ $\times$ $362\pm 20.00$
FedProx+PeWS $\mathbf{115\pm 5.29}$ $\mathbf{211\pm 16.52}$ $\mathbf{314\pm 27.83}$ $\mathbf{344\pm 27.30}$
Final accuracy after
$T$ collaboration rounds
FedProx+PeWS $\mathbf{99.96\pm 0.01}$ $\mathbf{98.40\pm 1.84}$ $\mathbf{90.40\pm 3.91}$ $\mathbf{99.92\pm 0.07}$
C.3 Wall-clock time vs. Accuracy
Figure 10 illustrates the wall-clock time versus accuracy results, which correspond to Figure 4 in the main paper. From this comparison, FedPeWS demonstrates a slightly improved performance over FedAvg in terms of wall-clock time in two of the four scenarios. However, it underperforms slightly in the remaining two scenarios, with only a marginal increase in time. This variance is attributed to the alternation between training masks and weights during the warmup phase, impacting the time efficiency.
Figure 10: Wall-clock time vs. accuracy plot corresponding to Figure 4 of the main paper. For better clarity, we selected a single run for each experiment at random.
C.4 FedPeWS-Fixed. Mask Length Study